# FaceFilterSense: A Filter-Resistant Face Recognition and Facial Attribute Analysis Framework

### v1

@@ 1.2 Motivations and Contribution
Few face filters may modify biometric features to a large extent and can promote the intent to impersonate or invade face recognition systems. The authors of Botezatu et al. botezatu used fun selfie filters in user authentication in their approach. The authors of Hedman et al. Hedman_2022 have tried to do similar work on Labelled Faces in the Wild Dataset (LFW). However, the variety of filters used in these approaches were limited and also the performance of existing state-of-the-art methods over filtered images was not up to the mark. To aid in counteracting the above-mentioned issues, we have developed a firstly created and published a public dataset of filtered images for 102 people having 2000+ images. We further developed a facial recognition system which performs well with filtered images as well.
@@ 2 Literature Review
Botezatu et al. botezatu $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
Hedman et al. Hedman_2022 $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
Rathgeb et al. Rathgeb $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
Serengil et al. serengil2020lightface serengil2021lightface $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
@@
    Fun Selfie Filters in Face Recognition: Impact Assessment and Removal: This work by Botezatu et al. botezatu very closely discusses the impact of fun selfie filters on facial recognition. However, most of the filters used by the authors are very basic occlusion filters like glasses, hats, bunny ears, etc. These filters introduce severe occlusion in the face which in turn decreases the recognition performance. However, it ignores the impact produced by filters that do not occlude the face but interfere with the facial key points and modify them to produce unidentifiable results. For e.g. Gender reverse or Child. Thus we were interested in including these filters. Also, the paper majorly focuses on the detection and removal of the filter.
@@
    On the Effect of Selfie Beautification Filters on Face Detection and Recognition: This paper by Hedman et al. Hedman_2022 uses a dataset of facial images applied with beautification and AR Filters on the standard LFW dataset. The filters applied by them include:
@@
    Figure 1: Examples of Filters used by Hedman et al.Hedman_2022
@@
    Differential Detection of Facial Retouching: A Multi-Biometric Approach: In this paper, Rathgeb et al. Rathgeb focused on the facial retouching techniques used by the filters in commonly used mobile applications like BeautyPlus, FotoRus, InstaBeauty etc. They have used multiple databases to test their results using models like COTS, SIFT, PRNU analysis, etc. The focus is majorly on the detection of the filter and filter-wise analysis, however, most of these apps provide beautification filters only.
@@
    LightFace: A Hybrid Deep Face Recognition Framework: This work by Serengil et al. serengil2020lightface provides a hybrid face recognition framework wrapping state-of-the-art models: VGG-Face, Google FaceNet, OpenFace, Facebook DeepFace, DeepID, ArcFace, Dlib, and SFace. They claim they have surpassed the human-level accuracy of 97.53% on facial recognition tasks. However facial filters are not considered in this work.
@@
    HyperExtended LightFace: A Facial Attribute Analysis Framework: This is another work by Serengil et al. serengil2021lightface provides a facial attribute analysis (age, gender, emotion, and race) framework for Python. They have analyzed how these parameters are related to facial features. However, they have also not worked around facial filters.
@@ 4.2 Performance Analysis and Comparison with existing approaches
The performance analysis of our custom models shows significant improvement over the existing state-of-the-art methods. We have used accuracy, precision, recall, and F-Score as the performance metric for the analysis of the classification models. While for the regression-based models, we used R2 Score, Mean Absolute Error, and Mean Squared Error as the performance metrics. For face recognition, we are comparing three state-of-the-art methods namely - ArcFaceDBLP:arcface , FaceNet512facenet and VGGFacevggface , provided by deepface library serengil2020lightface . ArcFace is a deep CNN-based model which is considered to be one of the most effective face recognition systems. Next, FaceNet comes with the inception module architecture which reduces the depth and number of trainable parameters without compromising on the accuracy. VGGFace is a face recognizer that performs significantly well in the wild. Table 2 shows the performance comparison of FaceFilterNet with these methods. This clearly shows how FaceFilterNet outperforms all these methods in the case of filtered images with an accuracy of 87.25%. This signifies that our model is able to recognize the correct person even after the application of filters.
@@
Performance Metric ArcFaceDBLP:arcface FaceNet512facenet VGGFacevggface FaceFilterNet
@@
Similarly, for the estimation of age, gender, and ethnicities from these filtered images, we used the HyperExtended LightFace serengil2021lightface which is a facial attribute analysis framework. It claims to have surpassed human-level accuracy with its deep CNN-based models. The experimental results of the models in this framework are reported in Table 3, 4, and 5 for Age, Gender, and Ethnicity respectively.
@@ 4.3.1 Filter-wise Analysis of Face Recognition
We consider a filter to be distorting if the average distance between the baseline image and the image filtered with that particular filter is very high. We calculated this metric for each filter to evaluate the average distortion produced by each filter. The results are shown in figure 5.
Figure 5: Amount of distortion produced by each filter to the face images in FRLL-Beautified
As per the standard threshold followed by Serengil et al., serengil2020lightface if the value of D given by equation 5 exceeds 0.75, it is considered to be breaking. We can clearly see from figure 5 that only the filter - Hipster Look from Snapchat is where our model fails. This can be seen as an indication that the amount of distortion produced by this filter is so high that the face becomes unrecognizable. Similarly, we can see the filters Sparkling Cartoon from Snapchat and Gender Reverse from FaceApp also produce a significantly large amount of distortion.
@@ 4.3.2 Filter-wise Analysis of Age Estimation
Figure 6: Filter-wise age mispredictions. The columns show the average reduction, increment, and net deviation in age estimation due to a given filter.
@@
Figure 6 shows the deviation in age estimation by the application of a filter that made the model estimate the age wrongly. It is clearly visible that in general, the model underestimates the age for each filter. This is also evident from the fact that the model produces a mean squared error of 1.74 years. However, filters like Hipster Look from Snapchat, Child Filter and Gender Reverse from FaceApp tend to make the faces look younger. Similarly, filters like Sparkling Cartoon from Snapchat make the faces look older.
@@ 4.3.3 Filter-wise Analysis of Gender Prediction
Figure 8: Filter-wise gender mispredictions: The columns show the number of instances the gender was wrongly predicted for a given filter
For the analysis of gender prediction, we tried to calculate the number of instances when the model predicted the wrong gender. Figure 7 shows filter-wise confusion matrices for gender predictions. Figure 8 shows the number of instances when the application of a filter made the model predict the gender wrongly. It is clearly visible that the number of times the model predicted the wrong gender (at max 5) is much lesser than the number of times it predicted correctly (at least 97) for a given filter. So the model is able to capture the artificial features introduced by filters to alter the gender. However, filters like - Gender Reverse from FaceApp produces some images which make them look more like a male. Similarly filters like - Hipster Look from Snapchat and Child Filter from FaceApp produce images that make them look more like a female. Thus these filters might be used maliciously to fake gender.
@@ 4.3.4 Filter-wise Analysis of Ethnicity Prediction
Figure 10: Filter-wise ethnicity mispredictions: Each column shows the number of instances when the wrongly predicted class was the given column name
For the analysis of ethnicity prediction, we tried to calculate the number of instances when the model predicted the wrong ethnicity. Figure 9 shows the filter-wise confusion matrix for ethnicity prediction. Figure 10 shows the number of instances when the application of a filter made the model predict the ethnicity wrongly. It is clearly visible that filters like - Child Filter from FaceApp and Puppy Filter from B612 misclassify most faces as east asian. This shows that filters like these might promote bias against this ethnicity. Similarly, Hipster Look from Snapchat misclassifies most faces as white. Thus these filters can be marked for scrutiny over ethnicity-based bias.

### v2

@@ 1.2 Motivations and Contribution
Few face filters may modify biometric features to a large extent and can promote the intent to impersonate or invade face recognition systems. The authors of Botezatu et al. [1] used fun selfie filters in user authentication in their approach. The authors of Hedman et al. [2] have tried to do similar work on Labelled Faces in the Wild Dataset (LFW). However, the variety of filters used in these approaches were limited and also the performance of existing state-of-the-art methods over filtered images was not up to the mark. To aid in counteracting the above-mentioned issues, we have developed a firstly created and published a public dataset of filtered images for 102 people having 2000+ images. We further developed a facial recognition system which performs well with filtered images as well.
@@ 2 Literature Review
Botezatu et al. [1] $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
Hedman et al. [2] $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
Rathgeb et al. [3] $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
Serengil et al. [4] [5] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
@@
    Fun Selfie Filters in Face Recognition: Impact Assessment and Removal: This work by Botezatu et al. [1] very closely discusses the impact of fun selfie filters on facial recognition. However, most of the filters used by the authors are very basic occlusion filters like glasses, hats, bunny ears, etc. These filters introduce severe occlusion in the face which in turn decreases the recognition performance. However, it ignores the impact produced by filters that do not occlude the face but interfere with the facial key points and modify them to produce unidentifiable results. For e.g. Gender reverse or Child. Thus we were interested in including these filters. Also, the paper majorly focuses on the detection and removal of the filter.
@@
    On the Effect of Selfie Beautification Filters on Face Detection and Recognition: This paper by Hedman et al. [2] uses a dataset of facial images applied with beautification and AR Filters on the standard LFW dataset. The filters applied by them include:
@@
    Figure 1: Examples of Filters used by Hedman et al.[2]
@@
    Differential Detection of Facial Retouching: A Multi-Biometric Approach: In this paper, Rathgeb et al. [3] focused on the facial retouching techniques used by the filters in commonly used mobile applications like BeautyPlus, FotoRus, InstaBeauty etc. They have used multiple databases to test their results using models like COTS, SIFT, PRNU analysis, etc. The focus is majorly on the detection of the filter and filter-wise analysis, however, most of these apps provide beautification filters only.
@@
    LightFace: A Hybrid Deep Face Recognition Framework: This work by Serengil et al. [4] provides a hybrid face recognition framework wrapping state-of-the-art models: VGG-Face, Google FaceNet, OpenFace, Facebook DeepFace, DeepID, ArcFace, Dlib, and SFace. They claim they have surpassed the human-level accuracy of 97.53% on facial recognition tasks. However facial filters are not considered in this work.
@@
    HyperExtended LightFace: A Facial Attribute Analysis Framework: This is another work by Serengil et al. [5] provides a facial attribute analysis (age, gender, emotion, and race) framework for Python. They have analyzed how these parameters are related to facial features. However, they have also not worked around facial filters.
@@ 4.2 Performance Analysis and Comparison with existing approaches
The performance analysis of our custom models shows significant improvement over the existing state-of-the-art methods. We have used accuracy, precision, recall, and F-Score as the performance metric for the analysis of the classification models. While for the regression-based models, we used R2 Score, Mean Absolute Error, and Mean Squared Error as the performance metrics. For face recognition, we are comparing three state-of-the-art methods namely - ArcFace[6], FaceNet512[7] and VGGFace[8], provided by deepface library [4]. ArcFace is a deep CNN-based model which is considered to be one of the most effective face recognition systems. Next, FaceNet comes with the inception module architecture which reduces the depth and number of trainable parameters without compromising on the accuracy. VGGFace is a face recognizer that performs significantly well in the wild. Table 2 shows the performance comparison of FaceFilterNet with these methods. This clearly shows how FaceFilterNet outperforms all these methods in the case of filtered images with an accuracy of 87.25%. This signifies that our model is able to recognize the correct person even after the application of filters.
@@
Performance Metric ArcFace[6] FaceNet512[7] VGGFace[8] FaceFilterNet
@@
Similarly, for the estimation of age, gender, and ethnicities from these filtered images, we used the HyperExtended LightFace [5] which is a facial attribute analysis framework. It claims to have surpassed human-level accuracy with its deep CNN-based models. The experimental results of the models in this framework are reported in Table 3, 4, and 5 for Age, Gender, and Ethnicity respectively.
@@ 4.3.1 Filter-wise Analysis of Face Recognition
We consider a filter to be distorting if the average distance between the baseline image and the image filtered with that particular filter is very high. We calculated this metric for each filter to evaluate the average distortion produced by each filter. The results are shown in Table 6.
Filter Name
Average L2-Euclidean
Distance
Haircut Filter FaceApp 0.458548
Child Filter FaceApp 0.507977
Gender Reverse Filter FaceApp 0.580298
Hipster Beard Style Filter FaceApp 0.332914
Hair Color Blonde Filter FaceApp 0.222967
Puppy Filter B612 0.494714
So Sad Filter B612 0.191711
Hipster Look Filter Snapchat 1.179643
Sparkling Cartoon Filter Snapchat 0.615396
Body Mellow Glow Filter Snapchat 0.427381
Table 6: Amount of distortion produced by each filter to the face images in FRLL-Beautified
As per the standard threshold followed by Serengil et al., [4] if the value of D given by equation 5 exceeds 0.75, it is considered to be breaking. We can clearly see from Table 6 that only the filter - Hipster Look from Snapchat is where our model fails. This can be seen as an indication that the amount of distortion produced by this filter is so high that the face becomes unrecognizable. Similarly, we can see the filters Sparkling Cartoon from Snapchat and Gender Reverse from FaceApp also produce a significantly large amount of distortion.
@@ 4.3.2 Filter-wise Analysis of Age Estimation
Filter Name
Average Age
Reduction (yrs)
Average Age
Increment (yrs)
Net Deviation (yrs)
Haircut Filter FaceApp -2.062500 1.478261 -0.292120
Child Filter FaceApp -3.371795 1.888889 -0.741453
Gender Reverse Filter FaceApp -3.492754 2.$0$0.746377
Hipster Beard Style Filter FaceApp -2.530303 1.812500 -0.358902
Hair Color Blonde Filter FaceApp -2.369863 1.636364 -0.366750
Puppy Filter B612 -3.200000 1.800000 -0.700000
So Sad Filter B612 -2.142857 1.733333 -0.204762
Hipster Look Filter Snapchat -5.774194 2.166667 -1.803763
Sparkling Cartoon Filter Snapchat -3.208333 3.153846 -0.027244
Body Mellow Glow Filter Snapchat -2.716049 1.200000 -0.758025
Table 7: Filter-wise age mispredictions: The columns show the average reduction, increment, and net deviation in age estimation due to a given filter.
@@
Table 7 shows the deviation in age estimation by the application of a filter that made the model estimate the age wrongly. It is clearly visible that in general, the model underestimates the age for each filter. This is also evident from the fact that the model produces a mean squared error of 1.74 years. However, filters like Hipster Look from Snapchat, Child Filter and Gender Reverse from FaceApp tend to make the faces look younger. Similarly, filters like Sparkling Cartoon from Snapchat make the faces look older.
@@ 4.3.3 Filter-wise Analysis of Gender Prediction
Filter
Number of Males
classified as Females
Number of Females
classified as Males
Haircut Filter FaceApp 1 0
Child Filter FaceApp 4 0
Gender Reverse Filter FaceApp 1 3
Hipster Beard Style Filter FaceApp 0 0
Hair Color Blonde Filter FaceApp 1 0
Puppy Filter B612 1 1
So Sad Filter B612 0 0
Hipster Look Filter Snapchat 5 2
Sparkling Cartoon Filter Snapchat 1 0
Body Mellow Glow Filter Snapchat 1 0
Table 8: Filter-wise gender mispredictions: The columns show the number of instances the gender was wrongly predicted for a given filter
For the analysis of gender prediction, we tried to calculate the number of instances when the model predicted the wrong gender. Figure 5 shows filter-wise confusion matrices for gender predictions. Table 8 shows the number of instances when the application of a filter made the model predict the gender wrongly. It is clearly visible that the number of times the model predicted the wrong gender (at max 5) is much lesser than the number of times it predicted correctly (at least 97) for a given filter. So the model is able to capture the artificial features introduced by filters to alter the gender. However, filters like - Gender Reverse from FaceApp produces some images which make them look more like a male. Similarly filters like - Hipster Look from Snapchat and Child Filter from FaceApp produce images that make them look more like a female. Thus these filters might be used maliciously to fake gender.
@@ 4.3.4 Filter-wise Analysis of Ethnicity Prediction
Filter black east asian west asian white
Haircut Filter FaceApp 0 3 6 6
Child Filter FaceApp 0 10 8 3
Gender Reverse Filter FaceApp 0 0 2 5
Hipster Beard Style Filter FaceApp 0 8 6 5
Hair Color Blonde Filter FaceApp 0 3 3 6
Puppy Filter B612 0 15 7 2
So Sad Filter B612 0 6 7 1
Hipster Look Filter Snapchat 0 3 0 29
Sparkling Cartoon Filter Snapchat 0 5 2 7
Body Mellow Glow Filter Snapchat 0 3 4 6
Table 9: Filter-wise ethnicity mispredictions: Each column shows the number of instances when the wrongly predicted class was the given column name.
For the analysis of ethnicity prediction, we tried to calculate the number of instances when the model predicted the wrong ethnicity. Figure 6 shows the filter-wise confusion matrix for ethnicity prediction. Table 9 shows the number of instances when the application of a filter made the model predict the ethnicity wrongly. It is clearly visible that filters like - Child Filter from FaceApp and Puppy Filter from B612 misclassify most faces as east asian. This shows that filters like these might promote bias against this ethnicity. Similarly, Hipster Look from Snapchat misclassifies most faces as white. Thus these filters can be marked for scrutiny over ethnicity-based bias.